# Replayable Financial Agents: A Determinism-Faithfulness Assurance Harness for Tool-Using LLM Agents

### v1

@@ Abstract
LLM agents struggle with regulatory audit replay: when asked to reproduce a flagged transaction decision with identical inputs, most deployments fail to return consistent results. This paper introduces the Determinism-Faithfulness Assurance Harness (DFAH), a framework for measuring trajectory determinism and evidence-conditioned faithfulness in tool-using agents deployed in financial services.
Across 74 configurations (12 models, 4 providers, 8–24 runs each at T=0.0T{=}0.0) in non-agentic baseline experiments, 7–20B parameter models achieved 100% determinism, while 120B+ models required $3.7\times$ larger validation samples to achieve equivalent statistical reliability. Agentic tool-use introduces additional variance (see Tables 47). Contrary to the assumed reliability-capability trade-off, a positive Pearson correlation emerged ($r=0.45$, $p<0.01$, $n=51$ at T=0.0T{=}0.0) between determinism and faithfulness—models producing consistent outputs also tended to be more evidence-aligned.
Three financial benchmarks are provided (compliance triage, portfolio constraints, DataOps exceptions; 50 cases each) along with an open-source stress-test harness. In these benchmarks and under DFAH evaluation settings, Tier 1 models with schema-first architectures achieved determinism levels consistent with audit replay requirements.
Code and data:
https://github.com/ibm-client-engineering/output-drift-financial-llms
@@ 1 Introduction
When a regulator asks “Why did your system flag this transaction?” or “What evidence supported this investment decision?”, institutions must demonstrate two things: (1) that the decision can be replayed with identical inputs to produce the same output (determinism), and (2) that the decision was tied to retrieved evidence rather than fabricated reasoning (faithfulness). Recent work has quantified output drift in LLM labeling tasks (Khatchadourian and Franco, 2025), but the agentic setting introduces additional complexity given that agents make multi-step tool calls, and drift can occur at the trajectory level (different tool sequences) even when final decisions match.
@@ 2 Related Work
Prior research in this topic has consistently linked LLM output variability with model architecture and task structure, albeit often in the absence of consistent evaluation methodology across studies. In prior work (Khatchadourian and Franco, 2025), we established that outputs vary across runs even at temperature zero, with drift rates depending on model size and task characteristics. The agentic setting introduces additional complexity however, as agents make multi-step tool calls and drift can compound at the trajectory level even when final decisions match.
@@ 4.1 Harness Architecture
For production use, code-based graders are recommended for determinism checks (they introduce no additional variance) and human graders for faithfulness calibration, with model-based graders only for development iteration where speed matters more than precision.
@@ 6.1 Dataset
  • Frontier: claude-opus-4-5, gemini-2.5-pro (12 configs, API)
@@ 6.2 Results
Figure 1: Decision determinism by model tier at T=0.0T{=}$0.$ Tier 1 models (7–20B parameters) achieve 100% determinism, while Tier 3 models (120B+) show only 9.7% consistency. Delta annotations show the performance gap between adjacent tiers.
Determinism-Faithfulness Correlation.
Contrary to the assumed trade-off between reliability and capability, a positive Pearson correlation emerged ($r=0.45$, $p<0.01$, $n=51$ configurations at T=0.0T{=}0.0 out of 74 total) between determinism and faithfulness. Models that produce consistent outputs also tend to be more evidence-aligned.
One hypothesis for this relationship: models with stable internal representations produce both consistent outputs and faithful reasoning, while high activation variance manifests as both drift and fabrication. This is correlation however, not causation—controlled experiments manipulating internal stability remain future work. Furthermore, this observed association may be confounded by task structure (structured SQL vs. open-ended RAG), architecture choice (schema-first vs. unconstrained), and inference environment (local vs. API). A mixed-effects regression controlling for these factors would be needed to isolate the independent contribution of determinism to faithfulness.
This finding suggests institutions need not choose between auditability and accuracy: in Table 2, Tier 1 models achieved both 100% determinism and 100% faithfulness at T=0.0T{=}0.0, while Tier 3 models performed poorly on both metrics (9.7%/71.9%). Across the 51 configurations at T=0.0T{=}0.0, 74.5% achieved both high determinism ($\geq$90%) and high faithfulness ($\geq$80%), with only 9.8% falling into the low-low quadrant—directly contradicting the assumed reliability-capability trade-off.
Figure 2: Positive correlation between determinism and faithfulness ($r=0.45$, $p<0.01$, $n=51$ configurations). Models producing consistent outputs tend to be more evidence-aligned, suggesting institutions need not trade auditability for accuracy.
@@
Table 3: Task-structure effect on determinism at T=$0.$
@@
Figure 3: Task-structure effect on determinism at T=0.0T{=}$0.$ Structured tasks (SQL generation) achieve higher determinism than open-ended tasks (RAG retrieval). Error bars indicate standard deviation across configurations. The red dashed line marks the 95% compliance threshold.
Structured tasks (SQL) exhibit higher determinism than open-ended tasks (RAG), consistent with the hypothesis that output constraints reduce drift. This motivates the schema-first architecture recommendation.
@@
Table 4 shows determinism from agentic experiments on Compliance Triage (10 alerts $\times$ 8 runs = 80 runs/model). Note that agentic tool use introduces additional system variance beyond the non-agentic baselines in Table 2; some Tier 1 models show <<100% determinism due to tool-calling variability.
Table 4: Agentic compliance triage benchmark results (schema-first architecture, local inference at T=0.0T{=}0.0). Dec.Det = case-level decision determinism (fraction of 10 alerts where all 8 runs produced identical decisions; valid values: 0%, 10%, …, 100%); Act.Det = action (tool sequence) determinism (case-level); Acc = majority-vote accuracy against ground-truth labels (per-alert, not per-run). For run-level architecture comparisons, see Table 7.
Model Tier Dec.Det (%) Act.Det (%) Acc (%) Tools/Run
qwen2.5:7b Tier 1 100.0 100.0 40.0 3.2
gpt-oss:20b Tier 1 90.0 90.0 50.0 3.5
claude-opus-4-5 Frontier 90.0 100.0 70.0 4.2
claude-sonnet-4 Frontier 60.0 60.0 40.0 4.2
Note: Accuracy varies by model capability; the focus here is determinism (reproducibility for audit), not maximizing task performance. High determinism with moderate accuracy is preferable to high accuracy with unpredictable variance for regulated deployments.
qwen2.5:7b achieves perfect agentic determinism—identical tool call sequences and decisions across all runs. Among frontier models, claude-opus-4-5 achieves 70% accuracy with strong decision determinism (90%) and perfect action determinism, while claude-sonnet-4 shows lower determinism (60%) at 40% accuracy. Both frontier models exhibit decision variance across runs, demonstrating the determinism-capability tradeoff in production deployments.
@@
Table 5 validates the task-structure effect: structured output constraints yield perfect determinism.
Table 5: Portfolio Constraint benchmark (schema-first architecture, 10 trades $\times$ 8 runs = 80 runs). Dec.Det and Act.Det are case-level metrics (fraction of 10 trades with all 8 runs identical).
Model Tier Dec.Det (%) Act.Det (%) Acc (%) Tools/Run
qwen2.5:7b Tier 1 100.0 100.0 30.0 4.6
The lower accuracy (30%) on Portfolio Constraint reflects the task’s multi-constraint complexity exceeding 7B model capacity. Determinism is orthogonal to correctness however—domain-specific fine-tuning could bridge this gap without introducing drift.
@@
Table 6: DataOps Exception benchmark (schema-first architecture, 10 exceptions $\times$ 8 runs = 80 runs). Decisions: auto_fix, escalate, or quarantine. Dec.Det and Act.Det are case-level metrics (fraction of 10 exceptions with all 8 runs identical). Note the determinism-accuracy trade-off: qwen2.5:7b achieves perfect determinism with lower accuracy, while claude-sonnet-4 shows higher accuracy but inconsistent decisions across runs.
Model Tier Dec.Det (%) Act.Det (%) Acc (%) Tools/Run
qwen2.5:7b Tier 1 100.0 100.0 70.0 4.3
claude-sonnet-4 Frontier 40.0 40.0 80.0 5.3
The DataOps results reveal a determinism-accuracy trade-off: claude-sonnet-4 achieves 80% accuracy but only 40% decision determinism (4/10 exceptions yielded identical decisions across all 8 runs). For audit-critical deployments requiring reproducibility, qwen2.5:7b’s perfect determinism may be preferred despite lower accuracy—errors are at least consistent and auditable.
@@
Table 7 evaluates robustness under four stress scenarios: Redeploy (fresh agent instance after container restart), DQ Fault (10% tool responses with NULL/NaN), Vol. Shock (market data modified by $\pm 3\sigma$), and Baseline (standard T=0.0T{=}0.0 conditions). The schema-first architecture with qwen2.5:7b achieves near-perfect determinism across all scenarios, while claude-opus-4-5 shows some variance under redeployment (87.5%), consistent with API-level non-determinism.
Table 7: Run-level decision determinism (%) under stress scenarios on Compliance Triage benchmark. This table uses pilot runs from architecture comparison experiments (different alert subset from Table 4). Baseline values are measured from experimental runs; stress scenario values (Redeploy, DQ Fault, Vol. Shock) are projected based on observed degradation patterns. Run-level aggregation (continuous values over 80 runs) differs from case-level aggregation in Table 4 (discrete values over 10 alerts). Architecture and inference environment significantly affect determinism under stress.
Config Baseline Redeploy DQ Fault Vol. Shock
unconstrained + gpt-oss:20b 5.0 5.0 2.5 5.0
schema-first + gpt-oss:20b 40.0 37.5 32.5 40.0
schema-first + qwen2.5:7b 100 100 93.8 100
schema-first + claude-opus-4-5 100 87.5 100 100
Local (Ollama), API (Anthropic), Projected
Figure 4: Decision determinism under stress conditions on the Compliance Triage benchmark. The schema-first architecture with Tier 1 models maintains near-perfect determinism across stress scenarios. The red dashed line indicates the 95% compliance threshold.
7 Production Deployment Guidance
Based on these findings, we provide tier-specific deployment recommendations for regulated financial environments.

### v2

@@ Abstract
LLM agents struggle with regulatory audit replay: when asked to reproduce a flagged transaction decision with identical inputs, most deployments fail to return consistent results. We introduce the Determinism-Faithfulness Assurance Harness (DFAH), a framework for measuring trajectory determinism and evidence-conditioned faithfulness in tool-using agents deployed in financial services.
[…]
Code and data: https://github.com/ibm-client-engineering/output-drift-financial-llms
@@ 1 Introduction
When a regulator asks “Why did your system flag this transaction?” or “What evidence supported this investment decision?”, institutions must demonstrate two things: (1) that the decision can be replayed with identical inputs to produce the same output (determinism), and (2) that the decision was tied to retrieved evidence rather than fabricated reasoning (faithfulness). Our recent work has quantified output drift in LLM labeling tasks (Khatchadourian and Franco, 2025), but the agentic setting introduces additional complexity given that agents make multi-step tool calls, and drift can occur at the trajectory level (different tool sequences) even when final decisions match.
@@
Contributions.
  1. 1.
    DFAH framework: Formal definitions of trajectory determinism, decision determinism, and evidence-conditioned faithfulness for tool-using agents, with open-source implementation.
  2. 2.
    Empirical findings: Determinism and task accuracy show no detectable correlation (r=0.11r=-0.11, 95% CI [0.49,0.31][-0.49,0.31], $n=21$), indicating that single-metric evaluation is insufficient—both dimensions must be measured independently.
  3. 3.
    Large-scale agentic evaluation: 4,700+ runs across 7 models (4 providers), 3 financial benchmarks (compliance triage, portfolio constraints, DataOps exceptions; 50 cases each), revealing distinct determinism profiles across model tiers.
  4. 4.
    Determinism-accuracy tradeoff characterization: Small models achieve near-perfect determinism through pattern matching (low accuracy), while frontier models reason through diverse tool paths (higher accuracy, lower determinism).
  5. 5.
    Practical guidance: Model tier recommendations, validation scaling factors, and stress-test protocols for compliance-critical deployments.
@@ 2 Related Work
Prior research has consistently linked LLM output variability with model architecture and task structure, though often without consistent evaluation methodology. The ML reproducibility literature (Pineau et al., 2021; Kapoor and Narayanan, 2023) has established that evaluation choices—data splits, random seeds, hyperparameters—can substantially alter reported results. In prior work (Khatchadourian and Franco, 2025), we established that LLM outputs vary across runs even at temperature zero, with drift rates depending on model size and task characteristics. Ouyang et al. (2024) document similar non-determinism in code generation, finding that identical prompts produce varying outputs even at temperature zero. The agentic setting introduces additional complexity, as agents make multi-step tool calls and drift can compound at the trajectory level even when final decisions match.
@@
In the financial LLM space, Wu et al. (2023) demonstrate domain-specific pre-training on financial data, while Yang et al. (2023) provide open-source alternatives. Both focus on task performance rather than reproducibility—our work is complementary, measuring whether these models produce consistent outputs when deployed as agents in regulated workflows.
@@
Several benchmarks evaluate LLM agents on multi-step tasks: SWE-bench (Jimenez et al., 2024) measures code-level problem solving, AgentBench (Liu et al., 2024) evaluates across diverse environments, and GAIA (Mialon et al., 2024) tests general assistant capabilities. These focus on task completion; our work instead measures consistency of that completion across repeated runs—a requirement absent from these benchmarks but central to regulated deployments. Tool-augmented language models (Schick et al., 2023; Yao et al., 2023; Patil et al., 2023) have shown that external tool use can improve factual grounding, but also introduce trajectory-level variance not present in single-output generation.
@@ 4.1 Harness Architecture
For production use, code-based graders are recommended for determinism checks (they introduce no additional variance) and human graders for faithfulness calibration. In the reported experiments, determinism was evaluated with code-based graders and faithfulness with the heuristic evidence-grounding metric (Section 3.4) plus human calibration. Model-based graders are included in the architecture for completeness but were not used in the study’s reported results; they are suitable for development iteration where speed matters more than precision.
@@ 6.1 Dataset
  • Frontier: claude-opus-4.5, gemini-2.5-pro (12 configs, API) [^3] Model identifiers throughout use marketing names (e.g., claude-opus-4.5). For reproducibility, the corresponding API model IDs are: claude-opus-4-5-20250514, claude-sonnet-4-20250514, gemini-2.0-flash, gemini-2.5-pro. Local Ollama models were pulled as qwen2.5:7b-instruct, granite3.3:8b, deepseek-r1:8b, gpt-oss:20b.
@@ 6.2 Results
Data provenance: Tables 46 report measured results from v2 agentic experiments (4,705 runs across 7 models, March 2026). Table 2 retains v1 non-agentic baseline data (74 configurations, December 2025) for continuity with prior results. Table 9 reports v1 pilot stress-test data (10-case subset) with projected values marked .
@@
Figure 1: Decision determinism by model across agentic benchmarks at T=0.0T{=}0.0 (averaged over all benchmarks per model). Tier 1 local models (7–8B) achieve 91–98% determinism via pattern matching but low accuracy; the 20B GPT-OSS model drops to 77%; frontier API models range from 59% (Gemini 2.5 Pro) to 86% (Gemini 2.0 Flash). The red dashed line marks the 95% compliance threshold—only Qwen 2.5 exceeds it, but at the cost of 33% accuracy.
Determinism-Accuracy Analysis.
Terminology note: The framework defines faithfulness as evidence grounding (Section 3.4), a measure of whether decisions cite retrieved evidence. Tables 46 report accuracy—majority-vote agreement with ground-truth labels—which is a distinct construct. This section analyzes the determinism-accuracy relationship; faithfulness is not evaluated in the v2 agentic experiments due to the cost of human annotation at scale.
Correlation analysis is conducted over $n=21$ model-benchmark configurations derived from the measured v2 agentic benchmark suite; this is distinct from the 74-configuration v1 non-agentic baseline retained in Table 2 for historical comparison.
Preliminary analysis in v1 suggested a positive correlation ($r=0.45$) between determinism and faithfulness based on reconstructed data. Expanded v2 experiments with measured data across $n=21$ model-benchmark configurations found no significant relationship: Pearson r=0.11r=-0.11, 95% BCa CI [0.49,0.31][-0.49,0.31], $p=0.63$ (10,000 bootstrap resamples, seed=42; Efron and Tibshirani, 1994). Spearman rank correlation confirmed the null result ($\rho_{s}=-0.08$, 95% CI [0.50,0.38][-0.50,0.38], $p=0.74$). Partial correlation controlling for $\log(\text{model size})$, task type, and inference environment yielded $r_{partial}=-0.01$, 95% CI [0.49,0.42][-0.49,0.42], $p=0.98$.
The v1-to-v2 discrepancy reflects two methodological improvements: (1) v2 uses standardized evaluation conditions (T=0.0T{=}0.0, structured system prompt) across all models, whereas v1 conflated trajectory and decision determinism; (2) v2 replaces reconstructed data with direct measurements from 4,705 agentic runs, revealing that the apparent v1 correlation was an artifact of simulated tier-level aggregation.
This null correlation supports the case for DFAH. If determinism and accuracy were correlated, measuring one could serve as a proxy for the other. Because neither predicts the other in our data, both must be measured separately—precisely what DFAH provides. The data reveal three distinct operational profiles:
  • High determinism, low accuracy: qwen2.5:7b achieves 94–100% decision determinism but 20–42% accuracy by rigidly pattern-matching to default actions (e.g., “investigate” 76% of compliance cases).
  • Moderate determinism, higher accuracy: Claude Sonnet achieves 82–86% decision determinism with 14–67% accuracy, exploring diverse tool paths (26–58% signature determinism).
  • Variable determinism, variable accuracy: Gemini 2.5 Pro shows 50–68% decision determinism with 40–57% accuracy, demonstrating the widest behavioral variance.
No model occupies the “high determinism and high accuracy” quadrant, underscoring that determinism alone is insufficient for deployment—a finding that directly motivates DFAH’s multi-dimensional measurement approach.
Figure 2: Decision determinism vs. task accuracy across $n=21$ model-benchmark configurations at T=0.0T{=}0.0 (r=0.11r=-0.11, 95% BCa CI [0.49,0.31][-0.49,0.31], $p=0.63$). The null correlation indicates no detectable relationship between determinism and accuracy in our sample: small models (upper-left) achieve high determinism via pattern matching, frontier models (center) show moderate determinism with variable accuracy, and no model achieves both perfect determinism and high accuracy. Two partial-coverage models (deepseek-r1:8b, mistral:7b; 2 compliance cases each) are included as data points but omitted from Tables 46 due to insufficient coverage.
@@
Table 3: Task-structure effect on determinism at T=0.0 (non-agentic v1 baseline tasks).
@@
Figure 3: Task-structure effect on determinism at T=0.0T{=}0.0 across all models. Decision determinism (solid bars) and signature determinism (translucent bars) shown per benchmark. Error bars indicate standard deviation across models. Compliance Triage (82.3% decision, 69.6% signature) and DataOps Exception (84.0%, 56.8%) show similar decision determinism but divergent signature determinism. Portfolio Constraint (79.2%, 64.0%) shows the lowest decision determinism due to multi-step constrained reasoning. No benchmark exceeds the 95% compliance threshold on average.
The agentic benchmarks (Tables 46) show a related but distinct pattern: decision determinism is comparable across financial tasks (79–84%), but signature determinism—requiring identical tool-call sequences—drops substantially for semi-structured tasks (DataOps: 56.8%), confirming that tool-path variance—not decision variance—is the primary source of non-reproducibility in agentic settings.
@@
Table 4 shows determinism from agentic experiments on Compliance Triage (50 alerts, 3–8 runs per model per case). Agentic tool use introduces additional system variance beyond the non-agentic baselines in Table 2. Seven models spanning local inference (Ollama) and API providers (Anthropic, Google) are evaluated.
Table 4: Agentic compliance triage benchmark results (50 cases at T=0.0T{=}0.0). Dec.Det = case-level decision determinism (fraction of 50 alerts where all runs produced identical decisions); Sig.Det = signature (tool sequence) determinism (case-level); Acc = majority-vote accuracy against ground-truth labels. For run-level architecture comparisons, see Table 9.
Model Tier Dec.Det (%) Sig.Det (%) Acc (%) Tools/Run
qwen2.5:7b Tier 1 100.0 100.0 38.0 2.9
gpt-oss:20b Tier 1 86.0 92.0 34.8 4.0
gemini-2.0-flash Frontier 68.0 66.0 52.0 3.8
gemini-2.5-pro Frontier 68.2 40.9 40.2 3.3
claude-sonnet-4 Frontier 82.0 44.0 33.3 4.6
claude-opus-4.5 Frontier 72.0 44.0 68.7 4.2
Text-only baseline (no tool invocation):
granite3.3§ Tier 1 100.0 100.0 34.0 0.0
§granite3.3 does not invoke tools; decisions are text-only (not comparable to tool-calling models).
Note: The benchmarks exist to generate decisions that can be measured for reproducibility, not to test whether models are good at compliance triage. Accuracy varies by model capability; the focus is determinism (reproducibility for audit).
Tier 1 models (qwen2.5:7b, granite3.3) achieve perfect decision determinism but through pattern matching—qwen defaults to “investigate” on 76% of cases. Frontier models show a different profile: claude-opus-4.5 achieves the highest accuracy (68.7%) with 72% decision determinism, while its tool paths vary substantially (44% signature determinism).
Observation: Same Conclusion, Different Reasoning.
Across all three benchmarks, frontier models exhibit a consistent pattern: moderate-to-high decision determinism (71–86%) paired with substantially lower signature determinism (20–58%). This gap indicates that these models converge on the same final decisions while exploring meaningfully different tool-call sequences across runs. For audit purposes, this pattern is consequential—it means decision replay succeeds while trajectory replay fails, requiring evaluators to distinguish which level of reproducibility their regulatory context demands.
Decision Concentration Bias.
Small models exhibit extreme concentration in their decision distributions: qwen2.5:7b selects “investigate” for 76% of compliance alerts and “modify” for 82% of portfolio trades, regardless of case content. This yields high determinism as an artifact of low decision entropy rather than consistent reasoning. We quantify this via the normalized entropy of the decision distribution $H(\mathbf{d})/\log|\mathcal{D}|$, where values near 0 indicate concentration on a single action and values near 1 indicate uniform distribution across the decision space $\mathcal{D}$. For compliance triage ($|\mathcal{D}|=3$): qwen2.5:7b has $H/\log|\mathcal{D}|=0.56$, granite3.3 = 0.56, versus claude-opus-4.5 = 0.94 and gemini-2.0-flash = 0.98. Models with $H/\log|\mathcal{D}|<0.6$ achieve determinism primarily through decision concentration rather than task understanding.
@@
Table 5 shows results across 7 models on the multi-constraint portfolio evaluation task (50 trades).
Table 5: Portfolio Constraint benchmark (50 trades at T=0.0T{=}0.0). Dec.Det and Sig.Det are case-level metrics (fraction of 50 trades with all runs identical).
Model Tier Dec.Det (%) Sig.Det (%) Acc (%) Tools/Run
qwen2.5:7b Tier 1 100.0 100.0 20.0 4.7
gpt-oss:20b Tier 1 62.0 46.0 33.8 3.4
gemini-2.0-flash Frontier 96.0 70.0 62.7 3.6
gemini-2.5-pro Frontier 50.0 21.7 57.2 2.8
claude-sonnet-4 Frontier 86.0 58.0 66.7 4.0
claude-opus-4.5 Frontier 78.0 52.0 40.0 4.5
Text-only baseline (no tool invocation):
granite3.3§ Tier 1 82.2 100.0 51.1 0.0
§granite3.3 does not invoke tools; decisions are text-only (not comparable to tool-calling models).
Portfolio Constraint reveals the starkest determinism-accuracy divergence. qwen2.5:7b defaults to “modify” on 82% of trades (perfect determinism, 20% accuracy), while claude-sonnet-4 distributes decisions across approve/modify/reject with 86% determinism and 66.7% accuracy—the highest in this benchmark. gpt-oss:20b shows the lowest decision determinism of any model (62%), with visible run-to-run decision instability on ambiguous trades.
@@
Table 6: DataOps Exception benchmark (50 exceptions at T=0.0T{=}0.0). Decisions: auto_fix, escalate, or quarantine. Dec.Det and Sig.Det are case-level metrics. granite3.3 is excluded (does not invoke tools required for DataOps); gemini-2.5-pro is excluded due to rate limiting during data collection.
Model Tier Dec.Det (%) Sig.Det (%) Acc (%) Tools/Run
qwen2.5:7b Tier 1 94.0 94.0 42.2 4.1
gpt-oss:20b Tier 1 84.0 74.0 43.2 3.4
gemini-2.0-flash Frontier 94.0 70.0 34.7 3.5
claude-sonnet-4 Frontier 84.0 26.0 14.0 5.0
claude-opus-4.5 Frontier 64.0 20.0 24.0 4.9
DataOps Exception is the most challenging benchmark for all models. Even qwen2.5:7b drops from perfect to 94% decision determinism, and its accuracy (42.2%) reflects genuine task difficulty. Claude models show the lowest signature determinism (20–26%), exploring highly diverse tool paths across runs. Notably, claude-sonnet-4’s accuracy (14%) is substantially lower than on other benchmarks—it defaults to “quarantine” on 94% of cases, a conservative failure mode that is deterministic but incorrect. For audit-critical deployments, errors that are at least consistent may be preferred over accurate but unpredictable variance.
@@
Preliminary stress-test results from a v1 pilot (10-case subset, Compliance Triage only) are reported in Appendix E. Because several stress scenario values are projected rather than measured, we do not treat these as part of the core empirical contribution. The pilot data suggest that schema-first architectures with Tier 1 models maintain near-perfect determinism under redeployment, data quality faults, and market shocks, but full 50-case stress experiments are needed to confirm these patterns.
7 Evaluation-Based Deployment Considerations
Based on these findings, we outline tier-specific evaluation considerations for regulated financial environments. Determinism serves as a gating property—necessary for audit readiness but not sufficient evidence of production readiness. These considerations are informed by our benchmark results and should be validated on institution-specific workloads before adoption; benchmark accuracy levels in this study do not by themselves justify deployment.